Subject Vitamin D
1 55.994
2 47.346
3 51.744
4 51.171
5 57.201
6 56.836
7 53.378
8 43.720
9 49.885
10 49.398
11 50.763
12 58.868
13 53.006
14 55.206
15 61.534
16 52.371
17 54.461
18 53.017
19 48.578
20 49.415
21 51.758
22 57.314
23 57.444
24 48.401
25 48.328
26 49.784
27 48.893
28 55.942
29 46.437
30 55.696
31 51.958
32 51.040
33 56.077
34 50.727
35 57.855
36 57.818
37 55.631
38 59.430
39 53.673
40 45.377
41 56.994

What is the estimated 95% confidence interval (CI) of the average blood vitamin D level of US landscapers in ng/mL?

Please note the following: 1) you might calculate a CI that is different from any of the multiple choice options listed below due to rounding differences, therefore select the closest match; 2) ensure you use either the large or small sample CI formula as appropriate; and 3) you may copy and paste the data into Excel to facilitate analysis.

Select one:

a. 51.7 to 54.2 ng/mL

b. 45.5 to 62.3 ng/mL

c. 57.6 to 61.0 ng/mL

d. 57.3 to 62.9 ng/mL
Answer #1

Mean of sample data Standard deviation 52.93827 4.175183

\alpha= 1-0.95= 0.05

z0.05/2 = 1.960

confidence interval =

Lower bound = 52.938 - 1.960 * 4.175 /\sqrt{41} = 51.66 = 51.7

Lower bound = 52.938 + 1.960 * 4.175 /\sqrt{41} = 54.216

So answer is a.